OpenStudy (anonymous):

im having trouble understanding this question, can someone help me? -The cost to insure jewellery is a fixed amount plus a percentage of the value of the jewellery. It costs $32 to insure $1000 worth of jewellery or $44.50 to insure $3500 worth of jewellery. What is the fixed amount to insure jewellery?

OpenStudy (unklerhaukus):

y=mx+b where y is cost of insurance x is cost of the jewelry b is the fixed amount

OpenStudy (anonymous):

im still not getting this, i could see this being a y-y1=m(x-x1), but im not sure if that would work, could you explain it a little more